# Optimization of Module Transferability in Single Image Super-Resolution: Universality Assessment and Cycle Residual Blocks

### v1

@@ 5 Conclusions
In this study, we propose the concept of “Universality” to quantify the transferability of modules. We introduce the UAE to measure module universality and design two optimized variants, CRB and DCRB, on the basis of RB. Analysis on their back propagation process reveals the underlying mechanisms behind these optimization behaviors. Extensive experiments across multiple SISR scenarios (natural scenes, remote sensing data, extreme industrial environments, and real-world applications) demonstrate that our approach outperforms several state-of-the-art models. Moreover, the proposed optimization strategy is applicable to other plug-and-play modules (e.g., RCAB), providing a principled framework for future architectural advances.

### v2

@@ 5 Conclusions
In this study, we propose the concept of “Universality” to quantify the transferability of modules. We introduce the UAE to measure module universality and design two optimized variants, CRB and DCRB, on the basis of RB. Analysis on their back propagation process reveals the underlying mechanisms behind these optimization behaviors. Extensive experiments across multiple SISR scenarios and other low-level tasks demonstrate that our approach outperforms several state-of-the-art models. Moreover, the proposed optimization strategy is applicable to other basic modules, providing a principled framework for future architectural advances.
Author Contributions
Haotong Cheng: Conceptualization, formal analysis, methodology, software, writing-original draft. Zhiqi Zhang: Investigation, validation, visualization. Hao Li and Xinshang Zhang: Resources, supervision, writing-review & editing.
Conflicts of Interest
The authors declare no conflicts of interest.
Data Availability Statement
This study uses publicly available datasets. The datasets for training are available via the following links: DIV2K(https://data.vision.ee.ethz.ch/cvl/DIV2K/), BelT(https://github.com/CUMT-AIPR-Lab/CUMT-AIPR-Lab). Datasets for evaluation are accessible in the following link: Set5, Set14, B100, Urban100(https://gitcode.com/Resource-Bundle-Collection/efef9/). Other Low-Level datasets could be found directly in their corresponding references. Experimental datas are provided within the article.